Hi, I just need help on doing the steps. I figured out it was 30, but do not remember how to do it. Thanks!
Assume that body temperatures of healthy adults are normally distributed with a mean of 98.2°F and a standard deviation of 0.62°F. Is P(98 < xbar < 98.4) greater for n =15 or
n = 30?

Answer by stanbon(75887)

Assume that body temperatures of healthy adults are normally distributed with a mean of 98.2°F and a standard deviation of 0.62°F.
Is P(98 < xbar < 98.4) greater for n =15 or n = 30?
--------------
Standard deviation for the distribution of x-bar is s/sqrt(n)
So std is smaller when n is larger.
----
But z-value = (x-u)/[s/sqrt(n)]
so z-value and n are directly related;
so z-value is larger as n is larger
-----
Therefore the probability is greater when n = 30.
=====================
Note:
z(98) = (98-98.2)/[0.62/sqrt(15)] = -1.2493
z(98) = (98-98.2)/[0.62/sqrt(30)] = -1.7668
------------
Cheers,
Stan H.